Show your work!!!

1.   Suppose you did an experiment of rolling a die 100 times, and you recorded each outcome. Then, you created the frequency and probability distribution table as Table 1.

                                                              Table 1

Outcome

(x)

Frequency

Probability P(x)

1

13

0.13

2

18

0.18

3

12

0.12

4

17

0.17

5

20

0.20

6

20

0.20

Compute the probability of following events using Table 1, and show the answers as decimals such as 0.25. (4 points)

Hint: If Event D = {1, 3}, then, P (D) = P (1) + P (3) = 0.13 + 0.12 = 0.25

  1. Event E = {2, 3}
    1. (E) =
  2. Event T ={1,2,3, 4}
    1. (T) =
  3. Event A ={1,2,3, 4,5,6}
    1. (A) =
  4. Event B = {}

(B) =
Answer #1

a) P(E) = P(2) + P(3) = 0.18 + 0.12 = 0.3

Therefore P(E) = 0.3

b) P(T) = P(1) + P(2) + P(3) + P(4) = 0.13 + 0.18 + 0.12 + 0.17 = 0.6

Therefore P(T) = 0.6

c) P(A) = P(1) + P(2) + P(3) + P(4) + P(5) + P(6) = 1

Therefore P(A) = 1

d) As B is empty, therefore P(B) = 0